OpenStudy (anonymous):

rewrite 9/20 and 11/25 so they have a common denominator

OpenStudy (anonymous):

Well you know both of the denominators (20 and 25) have a common factor of 100, so to get both fractions the same denominator of 100. \[\frac{ 9 }{ 20 }*(\frac{ 5 }{ 5 }) and \frac{ 11 }{ 25 }*(\frac{ 4 }{ 4 })\] Thus, you get \[\frac{ 45 }{ 100 } and \frac{ 44 }{ 100 }\] Hope it helps:)